OpenStudy (sgt_sauris):

What is the slope and y-intercept of the following equation: y = 1470 – 2x / 3. I think the slope is 2 and the y-intercept 3. Am I correct?

OpenStudy (anonymous):

y=mx+b so m is the slope and b it the y-intercept im pretty sure

OpenStudy (sgt_sauris):

Yes that is correct. But I am confused with which is which in my equation.

OpenStudy (sgt_sauris):

So do you know if I am correct @Ethanwestbrooks?

OpenStudy (anonymous):

what is the 1470-2?

OpenStudy (sgt_sauris):

1468

OpenStudy (sgt_sauris):

But I cannot do that since it is 1470 - 2x not 1470 - 2

OpenStudy (anonymous):

is 1470 =Y

OpenStudy (sgt_sauris):

? I do not get what you mean?

OpenStudy (anonymous):

is the equation y=mx+b equal to 1470=2x+3

OpenStudy (sgt_sauris):

The equation I listed is: y = 1470 – 2x / 3

OpenStudy (sgt_sauris):

So it would make a little bit of sense for 1470 to be my slope right?

OpenStudy (anonymous):

yeah it would make sense but 2 would make even more sense so i would go with 2

OpenStudy (sgt_sauris):

Since m = slope and b = y-intercept

OpenStudy (anonymous):

yeah

OpenStudy (sgt_sauris):

However... y=mx+b y = 1470 – 2x / 3 Wouldn't m = 1470 and x = 2 since the 2 is next to the x?

OpenStudy (anonymous):

yes but i think that on a graph its negative so i think if you knew what x was then you would substitute it and then subtract 1470

OpenStudy (sgt_sauris):

I guess you are right. So I should go with 2 as my slope and 3 as my y-intercept?

OpenStudy (anonymous):

so if x=2 it would be 2x2=4, 1470-4=1466

OpenStudy (anonymous):

yes

OpenStudy (sgt_sauris):

Ok... I guess we should hope I get a good grade xD Thank You! Let me send you your medal.

OpenStudy (anonymous):

Thank you! tell me what you get!

OpenStudy (sgt_sauris):

Did you get the medal? Its my first time sending one?

OpenStudy (anonymous):

yes i did! haha im new at this too

OpenStudy (sgt_sauris):

Ok good... I will pm you my grade when I get it if you are online... Thank You! :)

OpenStudy (anonymous):

welcome